Sol Summary – February 27th

Crew 189 Sol Summary Report 27 Feb 2018

Sol 9

Summary Title: The day of emergency procedures

Author: Victoria Da-Poian

Mission Status: All nominal

Sol Activity Summary: This morning, after a small issue of alarm (1 hour late on our usual wake up), we had a nice physical training (I was really motivated and did a really nice session!). After our breakfast, Laurent (our HSO) gave us a briefing about the emergency procedures. The afternoon was dedicated to use all this theory during EVA. Louis, Benoit, Gabriel, Laurent and I went outside the Hab to simulate all these emergency procedures. After this interesting EVA, we spent one hour doing the human factors experiment of Gabriel. (I think I am already addicted to this game, and playing with Laurent and Alex is really fun).

Look Ahead Plan: We were already all busy the first days of our second week. Tomorrow, we will all work on our experiments and will not go to EVA. At the end of this week, we will welcome a French journalist and everyone seems happy to see a new face. We will all work on our experiments and have our routine (physical training in the morning, human factors experiments, bread competition…)

Anomalies in work: Water leak in the Hab Lower deck

Weather: Nice and really warm

Crew Physical Status: Fine, still a bit tired from the first week

EVA: 7th

Reports to be file:

Commander

Journalist

Engineering

Greenhab Officer

EVA

Support Requested: None

Sol Summary – February 26th

Hello again Capcom,

Here is today’s sol summary:

Crew 189 Sol Summary Report 26 Feb 2018

Sol 8

Summary Title: Getting the rhythm

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
Today started by a quick EVA, to change LOAC’s battery, followed by an exploration of White Canyon. We spent this afternoon working on experiments and cleaning up the Hab as lower deck was becoming very dirty. The crew is now in good shape and the three of us recording our physical performance this morning scored new records in almost every discipline.
Only some of us lack of sleep, but except from this, everybody is doing great! The only thing worrying us is reduced food’s portions, but we start to adapt to lighter meals.

Look Ahead Plan: Emergency procedures training in the morning, in situ testing during a quick EVA near the hab in the afternoon.

Anomalies in work: Water leak in the Hab Lower deck

Weather: Mostly sunny.

Crew Physical Status: Fine.

EVA: –

Reports to be file:
Commander
Journalist
Engineering
Greenhab Officer
EVA report
HSO report

Support Requested: –

Louis Mangin
Crew 189 Commander

Sol Summary – February 25th

Crew 189 Sol Summary Report 25 Feb 2018

Sol 7

Summary Title: Rest Day

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
We decided to declare two rest days this year because all the lack of
sleep we accumulated last year. No fixed hour of wake up for today,
but everybody was busy by 11:00 a.m. This day was supposed to be free
and dedicated to long run tasks or personal time. In the afternoon, we
split in different team, working in the greenhab to rearrange it,
cooking, producing an encouragement video for school’s student
selected for a TV show. It was also the time for a first individual
debriefing about the sport sessions I organize.

Look Ahead Plan: Change of LOAC’s battery, exploration of White Canyon
in EVA. Experiments repair, backpack change of battery, emergency
procedures inside training.

Anomalies in work: Water leak in the Hab Lower deck

Weather: Mostly sunny, cold.

Crew Physical Status: Fine.

EVA: –

Reports to be file:
Commander
Journalist
Engineering
Greenhab Officer

Support Requested: –

Sol Summary – February 21st

Crew 189 Sol Summary Report 21 Feb 2018

Sol 3

Summary Title: First results

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
Normal start: everybody is now accustomed to the physical training even if some legs are already heavy. The third EVA was supposed to deploy the LOAC but it didn’t go that well. The power system had an issue and wires melted once I pluged it in. We then brought the equipement back to the hab, and went on for the EVA, going East to do a bit of exploration near Cactus Road, and shoot photos for the school.
This afternoon was dense: after lunch break, we answered to a French journalist interview, played the first session of Gabriel’s human factor experiment, installed a new shade on the GreenHab, worked on experiments, so that we didn’t have enough time to fix the backpack’s fans. Tomorrow EVA will then be done using the new version, before we take time to take care of these (this time, I really want to do it tomorrow).

Look Ahead Plan: Third deployment try of LOAC, data gathering of MegaARES, exploration of Candor Chasma, backpack’s fans inspection, end of shade replacement in greenhab.

Anomalies in work: Broken light in the bathroom (no replacement one found). Water leak in the Hab Lower deck, broken fans in spacesuits.

Weather: Cold and Cloudy

Crew Physical Status: Fine, already a bit tired, but everybody naped.

EVA: Second

Reports to be filed:
Commander
Journalist
HSO
Engineering
Greenhab Officer
EVA

Support Requested: New light for bathroom (see Engineering report)

Sol Summary – February 20th

Crew 189 Sol Summary Report 20 Feb 2018

Sol 2

Summary Title: First experiments deployment

Author: Victoria Da-Poian

Mission Status: All nominal

Sol Activity Summary: Our day began with our lovely physical training.
It is already adopted by each crewmember. Since this morning, everyone
has experienced the EVAs. We managed to deploy one experiment (the
MegaARES) but a small problem enabled the LOAC deployment. We will fix
it with Jérémy in order to try again tomorrow. Even if Gabriel had a
very foggy helmet (fans problems), and everyone seems really happy and
I have to say the photos Benoit took are really nice. During our EVA,
the rest of the crew fixed the water boiler, cleaned, cooked and
worked on experiments. After the lunch and a little break, Gabriel
showed us his humans factors experiment, and everyone worked on their
own experiments and reports.

Look Ahead Plan: After our first deployment, the second air sensor,
(LOAC) will normally be deployed tomorrow.

Anomalies in work: Water leak in the Hab Lower deck, broken fans in spacesuits

Weather: Nice

Crew Physical Status: Fine, already a bit tired from the first EVAs

Reports to be filed
:
Commander
Journalist
HSO
Engineering
Greenhab Officer
EVA

Support Requested: No more boiler is needed. Would it be possible to
have the RoverCal that is in the RAM delivered in the airlock please?

Sol Summary – February 19th

Crew 189 Sol Summary Report 19 Feb 2018

Sol 1

Summary Title: The day of Firsts

Author: Victoria Da-Poian

Mission Status: All nominal

Sol Activity Summary: First physical training, first EVA. We started our day by the very famous physical training of the ISAE Supaero crews. After a nice morning working on our mission planning, fixing our experiments, and getting ready for our first EVA, we had a nice lunch with our first dried vegetables. After the lunch, the all crew worked on the EVA preparation to help the ones going outside and to learn the EVAs functioning. The EVA went fine and we will deploy our experiments tomorrow. Our first bread is baking.

Look Ahead Plan: Because of the wind planned for today, only one part of the crew went out for an EVA in the Hab surroundings to select a location to deploy our air sensors, MegaARES and LOAC. Tomorrow the other part of the crew will go to their first EVA to deploy our experiments.

Anomalies in work: Water leak in the Hab Lower deck, broken fans in spacesuits, water boiler broken

Weather: Windy, cold in the station

Crew Physical Status: Fine, already a bit tired from the first physical training.

EVA: First one

Reports to be file:

Commander

Journalist

HSO

Engineering

Greenhab Officer

Astronomer Officer

EVA

Support Requested: new water boiler (please)

Sol Summary – February 18th

Crew 189 Sol Summary Report 18 Feb 2018

Sol 0

Summary Title: Getting Ready

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary: Last briefings, beginning of the simulation. We
started by a round trip to Grand Junction to get back the rental car
while the rest of the crew was settling in the hab, learned about the
surroundings and station management with Dr Rupert. Then we learned
about driving, spacesuits, and had a briefing about simulation with Dr
Rupert. We then closed the airlock at 5:50 PM before sunset.

Look Ahead Plan: Because of the wind planned for tomorrow, we changed
of plans, and will only go out for an EVA in the Hab surroundings to
select a location to deploy our air sensors, MegaARES and LOAC.

Anomalies in work: Water leak in the Hab Lower deck, broken fans in
spacesuits, water heater badly calibrated. Two greenhab thermometers
probably broken. Voltmeter not working.

Weather: Very windy, hard to see and to travel without space suits.

Crew Physical Status: Fine, already a bit tired from travel.

EVA: None

Reports to be file:
Commander
Journalist
HSO checklist
Engineering
Greenhab Officer

Support Requested: One or two thermometers.

Sol Summary Report – February 16th

Crew IKEA Sol Summary 16 Feb 2018

SOL 6:

Summary Title: Finally… Adventure Time

Author’s name: Atila Meszaros

Mission Status: All nominal

Sol Activity Summary:

I feel like I haven’t been in the Hab all day. The weather looked friendly and allowed us to make our first day of expedition. We spent most of the morning doing the growth mediums for our extremophiles, defining the microbiological points of interest and planning the activities for the arrival of the next crew.

The contractor already has almost the entire ladder of the lower deck made. Tomorrow they will come to finish it. It looks amazing. Tomorrow we will be sending some pics! While the guys worked, we managed to turn on Shannon’s car, apparently, it will not need a new battery. However, we shouldn’t take your eyes off and go carefully in case this event repeats. We picked up a package that contains equipment to record some footage at MDRS and we filled the aquaponics with water to have it almost ready by the time the next crew arrives. By the way, the Green Hab plants are looking amazing.

On our first day of adventure, we entered the Valley of the Stars in search of gypsum that may contain halophiles. Looking in any direction was simply spectacular. It is one of the most majestic places I’ve ever seen. We collected 5 different samples of vein, soil, and eruption. In a few days, we will be sending a Science Report with the findings of our first search.

It has been a very hectic day. But I am happy with everything we have achieved today. Now, I have to clean the Hab for the next crew and move to RAM (I’m going to be the first person to live there, I’m kind of excited).

Thanks for joining us tonight, Veronica!

How are you doing?

Atila

Look Ahead Plan: Recieve Supaero crew. I’m looking forward to meeting them!

Anomalies in work: None

Weather: Temperature: °6C

Crew Physical Status: Healthy

Reports to be filed:

Sol Summary

Operations Report

Science/Intern Report

Support Requested: None

Sol Summary – February 15th

Crew IKEA Sol Summary 15 Feb 2018

SOL 5:

Summary Title: Shelter from the Storm

Author’s name: Atila Meszaros

Mission Status: All nominal

Sol Activity Summary:

Another slow and very windy day. The ladders installations and the mud have kept us inside most of the day. However, we managed to go out and we drive Opportunity, Spirit, and Deimos. Due to the new ladder installation, I’ve been trapped in the upper deck almost all day and we had to postpone our adventures until (I hope) tomorrow. All the Hab is screaming because of the winds. It sounds pretty awesome. but I hope I can get some sleep

Thank you very much for being with us tonight!

Atila

Look Ahead Plan:
Getting a new battery for Shannon’s car, taking Rovers out and I’m rooting for finally some field work and exploration!

Anomalies in work:
Installation of brand new ladders for the lower deck

Weather: Temperature: 5°C

Crew Physical Status: Healthy

Reports to be filed:

Sol Summary
Operations Report

Support Requested: None

Sol Summary – February 14th

Crew IKEA Sol Summary 14 Feb 2018

SOL 4:

Summary Title: Here comes the sun
Author’s name: Atila Meszaros

Mission Status: All nominal

Sol Activity Summary:

Busy day for everyone. The contractor came and installed the new ladders for the upper deck and just before the end of the week the new lower deck ladder is going to be installed. They look fantastic! I’ll love to be here whit the Ikea project will be complete.

Sadly, the mud persisted due to the lack of wind and warmer weather, so we couldn’t get out and do some field work, tomorrow I bet will be luckier. Instead of going out, I worked all day on the Green Hab on the aquaponics, doing some harvest and organizing everything.

Everything is looking good on Mars, but the wind is sometimes scary, inside the Green Hab, you can almost feel the wind screaming. It’s pretty amazing, though.

Thank you for being with us tonight, Sylvain!

Atila

Look Ahead Plan:

Getting a new battery for Shannon’s car, taking Rovers out and I’m rooting for finally some field work and exploration!

Anomalies in work:

Installation of brand new ladders for the upper deck
Shannon’s car battery died

Weather: Temperature: 6°C

Crew Physical Status: Healthy

Reports to be filed:

Sol Summary
Operations Report

Support Requested: None